OpenStudy (xxlxxlxxlxlxlxlx):

The probability of event A is x, and the probaility of event B is y. If the two events are independent, which condition must be true? P(A|B) = x P(A|B) = y P(B|A) = x P(B|A) = xy

OpenStudy (holsteremission):

If \(A\) and \(B\) are independent, then \(\mathbb{P}(A|B)=\dfrac{\mathbb{P}(A\cap B)}{\mathbb{P}(B)}=\dfrac{\mathbb{P}(A)\mathbb{P}(B)}{\mathbb{P}(B)}=\mathbb{P}(A)\). Similarly, \(\mathbb{P}(B|A)=\mathbb{P}(B)\).